# A note on Majorana representation of quantum states

### v3

@@ 1 Introduction
In this note, we establish a connection between the Majorana representation and symmetry class of tensors in $\mathbb{V}^{\otimes(d-1)}$ for $\mathbb{V}=\mathbb{C}^{2}$ associated with the principal character $\xi$. Using this connection, we provide an easy scheme to determine $v_{1},\dots,v_{d-1}\in\mathbb{C}^{2}$ associated with a given vector $(a_{0},\dots,a_{d-1})\in\mathbb{C}^{d}$. Additionally, we present a simple formula for the inner product of $\mathbf{a}$ and $\mathbf{b}$ in $\mathbb{C}^{d}$ in terms of their $(d-1)$-qubit presentations. Numerical examples are given to illustrate the result. Extension of the result to mixed states is also considered.
@@ 2.1 Preliminary
$T(v_{1}\otimes\cdots\otimes v_{d-1})=\frac{1}{(d-1)!}\sum_{\sigma\in S_{d-1}}\xi(\sigma)v_{\sigma^{-1}(v_{1})}\otimes\cdots\otimes v_{\sigma^{-1}(d-1)}.$
Then $\mathbb{V}_{\xi}^{(d-1)}=T(\mathbb{V}^{\otimes(d-1)})$ is a subspace of $\mathbb{V}^{\otimes(d-1)}$ known as the symmetry class of tensors over $\mathbb{V}$ associated with $\xi$ on $S_{d-1}$. The elements in $\mathbb{V}_{\xi}^{(d-1)}$ of the form $T(v_{1}\otimes\cdots\otimes v_{m})$ are called decomposable tensors and are denoted by $v^{\bullet}=v_{1}\bullet\cdots\bullet v_{m}$. One may see [2, 4] for some general background. In fact, researchers have used decomposable tensors to model boson states; see [1].
@@
If $v_{1},\dots,v_{d-1}\in\mathbb{C}^{2}$, then the decomposable tensor
@@ 2.2 Main result and examples
By Theorem 2.1, every vector $f^{\bullet}\in\mathbb{V}_{\xi}^{(d-1)}$ admits a representation of the form $u_{1}\bullet\cdots\bullet u_{d-1}$. In particular, if $(a_{0},\dots,a_{d-1})^{t}\in\mathbb{C}^{d}$ is a quantum sate, i.e., a unit vector, then
$a_{0}f_{0}^{\bullet}+\cdots+a_{d-1}f_{d-1}^{\bullet}=\frac{1}{\gamma}\left(\frac{v_{1}}{\|v_{1}\|}\bullet\cdots\bullet\frac{v_{d-1}}{\|v_{d-1}\|}\right),$
where $v_{1},\dots,v_{d-1}$ are defined as in Theorem 2.1 and
@@ 2.3 Mixed states
Recall that a general quantum state is called a mixed state and is represented by a density matrix $\rho$, which is a positive semi-definite matrix with trace 1. If $\rho={\bf a}{\bf a}^{*}$ for a unit vector ${\bf a}=(a_{0},\dots,a_{d-1})^{t}\in\mathbb{C}^{d}$, then $\rho$ is pure sate. If ${\bf a}$ corresponds to $u_{1}\bullet\cdots\bullet u_{d-1}\in\mathbb{V}_{\xi}^{d-1}$, then by (3) the corresponding density matrix $\rho={\bf a}{\bf a}^{*}\in\mathbb{M}_{d}$ has $(r,s)$ entry equal to
@@
Alternatively, by purification one may express $\rho$ as the partial trace of a pure stat $|\psi{\rangle}{\langle}\psi|$ with $|\psi{\rangle}\in\mathbb{C}^{d^{2}}$, which admits a Majorana representation of $d^{2}-1$ qubits.

### v4

@@ 1 Introduction
In this note, we establish a connection between the Majorana representation and the symmetry class of tensors in $\mathbb{V}^{\otimes(d-1)}$ for $\mathbb{V}=\mathbb{C}^{2}$ associated with the principal character $\xi$. Using this connection, we provide an easy scheme to determine $v_{1},\dots,v_{d-1}\in\mathbb{C}^{2}$ associated with a given vector $(a_{0},\dots,a_{d-1})\in\mathbb{C}^{d}$. Additionally, we present a simple formula for the inner product of $\mathbf{a}$ and $\mathbf{b}$ in $\mathbb{C}^{d}$ in terms of their $(d-1)$-qubit presentations. Numerical examples are given to illustrate the result. Extension of the result to mixed states is also considered.
@@ 2.1 Preliminary
$T(v_{1}\otimes\cdots\otimes v_{d-1})=\frac{1}{(d-1)!}\sum_{\sigma\in S_{d-1}}\xi(\sigma)v_{\sigma^{-1}(1)}\otimes\cdots\otimes v_{\sigma^{-1}(d-1)}.$
Then $\mathbb{V}_{\xi}^{(d-1)}=T(\mathbb{V}^{\otimes(d-1)})$ is a subspace of $\mathbb{V}^{\otimes(d-1)}$ known as the symmetry class of tensors over $\mathbb{V}$ associated with $\xi$ on $S_{d-1}$. The elements in $\mathbb{V}_{\xi}^{(d-1)}$ of the form $T(v_{1}\otimes\cdots\otimes v_{d-1})$ are called decomposable tensors and are denoted by $v^{\bullet}=v_{1}\bullet\cdots\bullet v_{d-1}$. One may see [2, 4] for some general background. In fact, researchers have used decomposable tensors to model boson states; see [1].
@@
Suppose $v_{1},\dots,v_{d-1}\in\mathbb{C}^{2}$. Then the decomposable tensor
@@ 2.2 Main result and examples
By Theorem 2.1, every vector $f^{\bullet}\in\mathbb{V}_{\xi}^{(d-1)}$ admits a representation of the form $u_{1}\bullet\cdots\bullet u_{d-1}$. In particular, if $(a_{0},\dots,a_{d-1})^{t}\in\mathbb{C}^{d}$ is a quantum state, i.e., a unit vector, then
$a_{0}f_{0}^{\bullet}+\cdots+a_{d-1}f_{d-1}^{\bullet}=\frac{\gamma_{r}}{|\gamma_{r}|\gamma}\left(\frac{v_{1}}{\|v_{1}\|}\bullet\cdots\bullet\frac{v_{d-1}}{\|v_{d-1}\|}\right),$
where $\gamma_{r},v_{1},\dots,v_{d-1}$ are defined as in Theorem 2.1 and
@@ 2.3 Mixed states
Recall that a general quantum state is called a mixed state and is represented by a density matrix $\rho$, which is a positive semi-definite matrix with trace 1. If $\rho$ is rank one, i.e., $\rho={\bf a}{\bf a}^{*}$ for a unit vector ${\bf a}=(a_{0},\dots,a_{d-1})^{t}\in\mathbb{C}^{d}$, then $\rho$ is pure state. If ${\bf a}$ corresponds to $u_{1}\bullet\cdots\bullet u_{d-1}\in\mathbb{V}_{\xi}^{d-1}$, then by (3) the corresponding density matrix $\rho={\bf a}{\bf a}^{*}\in\mathbb{M}_{d}$ has $(r,s)$ entry equal to
@@
Alternatively, by purification one may express $\rho$ as the partial trace of a pure stat $|\psi{\rangle}{\langle}\psi|$ with $|\psi{\rangle}\in\mathbb{C}^{rd}$, which admits a Majorana representation of $rd-1$ qubits.